OpenStudy (anonymous):

What is the slope of the line given by the equation below? y + 2 = -3(x - 5)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

is it -3

jimthompson5910 (jim_thompson5910):

yes because it's in the form y - y1 = m(x-x1) m = slope

jimthompson5910 (jim_thompson5910):

y - y1 = m(x-x1) is point slope form m = slope (x1,y1) = point